4x(x-6)-5x(x-5)=2 sovle for x. I must admit on this one I'm totaly lost.
Answer by scott8148(6628)

distributing ___ 4x^2 - 24x - 5x^2 + 25x = 2

collecting terms ___ -x^2 + x = 2 ___ 0 = x^2 - x + 2

use quadratic formula to find x